# Log-normal Superstatistics in the Confined Motion of Ants

### v4

@@ Abstract
Panic regime was examined by tracking 48 trajectories of ants during the next 10 minutes following repellent injection. Several observables remained unchanged (Fig. 8): mean acceleration stayed linear in speed, $\tau_{2}$ remained stable, and both $\gamma^{-1}$ and $D$ showed no significant changes. Centered acceleration distributions also retained their shape, indicating preserved local Gaussian equilibria for $t<\tau_{2}$. However, the velocity distributions developed stretched tails, consistent with gradual clustering in the safe cell, and non-gaussianity persisted due to the interplay of $v$-dependent noise and Superstatistics. The only clearly altered observable was $\sigma(v)$, which maintained isotropy but evolved over time. Specifically, two parameters, $\Delta\Sigma(t)$ and $v_{\sigma}(t)$, underwent a sharp, brief increase at minute $9.2$, then relaxed, while $\Sigma(t)$ stayed constant (Fig. 4a). For later simulations, $\Delta\Sigma(t)$ and $v_{\sigma}(t)$ were fit to contiguous back-to-back exponential functions with matching growth and relaxation rates, selected for their simplicity in characterizing rapid changes. The timing at which this peak occurs will subsequently be used as a key measure of one of the most significant timescales in the ants’ mechanism in response to panic.
@@
$\begin{split}\frac{d}{dt}v_{i}(t)&=-\gamma(v_{i}(t)-\tilde{v}_{i}(t))\\ &\quad\,+\sigma(v,t)\sqrt{D(t)}\,\xi_{i}(t)+F_{w}(r_{i})\end{split}$ (5a)
@@
Figure 4: (a) Time evolution of the parameters in Eq. (2). $\Delta\Sigma(t)$ and $v_{\sigma}(t)$ show a sharp change at minute 9.2, while $\Sigma(t)$ remains stable. (b) Probability of finding an ant from the safe cell at a distance less than or equal to $2\,\mathrm{cm}$ from the barrier. The last 5 minutes of the experiment are highlighted in orange, where the curve remains approximately constant near zero. During this interval, the ant cluster is stable.
@@
In summary, we have shown that Log-normal Superstatistics emerges as a robust statistical framework that captures the transition from normal to danger behavior in confined ants, while maintaining continuity and providing a foundation upon which panic-induced dynamics can be superimposed. We reported non-Gaussian behavior, a slowly varying diffusion coefficient, and a collective escape response triggered by inter-ant communication. A data-driven superstatistical model with Log-normal diffusion successfully reproduced these features and captured the danger transition, relying on a single free parameter. The model uncovers a fast, memory-limited information cascade followed by a slower collective reorganization into a stable cluster.
@@
Acknowledgements from A. Reyes to the Abdus Salam ICTP STEP program for financial support. We thank A. Celani, D. Gordon, R. Mulet, T. Tohme and D. Machado for valuable discussions, G. Simon for the 3D-printed of the cells, and A. Serrano for assistance with Fig. 1. E. Altshuler acknowledges an Invited Professorship supported by CNRS at the University of Lyon.
Appendices
@@
Figure 5: (a) Acceleration as a function of position (component wise). Inset: acceleration absolute value as a function of the distance from the walls. A characteristic length $\ell_{w}=0.6\,\mathrm{cm}$ is defined as the effective wall thickness. (b) The grey sections have been removed from a sample trajectory, as they are within a distance $\ell_{w}$ from the walls: in order to ensure that statistics is not contaminated by boundary effects, we only use for analysis the non-gray sections.
@@
I. Boundary effects: The baseline regime was used to evaluate boundary effects while measuring the extent of wall “forces”. Experimental acceleration components (i.e., $a_{x}$ and $a_{y}$) were arranged based on increasing values of $x$- and $y$- position components, respectively. Both acceleration time series were binned and a running average was taken (main panel of Fig. 5a). The ants slow down as they approach the walls, revealing the influence of the boundaries as if they were confined by a soft, square-box-like potential. Acceleration absolute value against the distance from the wall revealed a characteristic length $\ell_{w}=0.6\,\mathrm{cm}$, obtained by fitting an exponential function (inset of Fig. 5a). Henceforth, analysis was conducted with trajectory segments that were consistently at least $\ell_{w}$ units away from every boundary, to discard boundary effects (Fig. 5b).
II. Mechanisms for danger transmission: To isolate potential mechanisms of danger transmission, we used three barrier types (Fig. 6b): a transparent Glass wall allowing only visual cues; a Thick barrier restricting the interaction to mutual antennation; and a Thin barrier allowing full contact between ants in adjacent arenas (Fig. 6c).
@@
Figure 6: (a) Ant used in experiments: leafcutter ant Atta insularis, a species native to Cuba known as “bibijagua”. (b) Sketch of the three types of barriers used for the protocols Glass, Thick and Thin. The latter is also used for the Ctrl protocol. (c) Top view of ants near a thick and thin barriers. Short distance contacts are not allowed by the thick barrier. (d) Lateral view of an ant near a thin barrier. As a reference, the length of the ants bodies is around $1\,\mathrm{cm}$.
@@
To characterize collective motion across protocols, we tracked the ants’ CM in both arenas (Fig. 7a-d). In all cases, the CM in the danger cell remained near the geometric center regardless of barrier type. The same held for the safe cell, except under the Thin protocol, where the CM systematically shifted away from the barrier after repellent was introduced (Fig. 7c). Fig. 7e-h shows the time evolution of the horizontal CM position in the safe cell, averaged over 10 experimental repetitions. A clear displacement emerges only in the Thin case after the injection at minute 5. In contrast, the Ctrl case (i.e., where the danger cell is empty) shows no comparable motion, indicating that passive diffusion alone cannot account for the observed shift.
Figure 7: Snapshots of experiments (a) Glass, (b) Thick, (c) Thin and (d) Ctrl. Green, blue, and red traces show the center of mass (CM) trajectory for consecutive 5-minute windows. (e-h) Horizontal CM evolution in the safe cell, relative to its initial position. Shading shows standard deviation. (i) Net horizontal displacement of the ants CM based on 10 repetitions of each type of experiment.
@@
Figure 8: (a) Speed dependence of parallel and perpendicular acceleration components across time intervals. (b) Local flatness vs lag time, computed over trajectory segments from time intervals $[0,5)$ to $[14,15)$ minutes. Final time of each interval is shown in legend. (c) Velocity autocorrelation functions (VACF) vs lag time. The solid line has slope $\gamma=(0.53\,\mathrm{s})^{-1}$. Notice that the slow time scale $\tau_{2}$ is not captured by the VACF. (d) Probability density function (PDF) of centered parallel and perpendicular components of acceleration fluctuation. (e) PDF of Cartesian velocity components. (f) Temporal and speed dependence of the intensity of acceleration fluctuations for the parallel (blue) and perpendicular (orange) components after panic was induced. Continuous lines show fits to Eq. (2).

### v5

@@ IV Panic regime
Several observables remained unchanged (Fig. 8): mean acceleration stayed linear in speed, $\tau_{2}$ remained stable, and both $\gamma^{-1}$ and $D$ showed no significant changes. Centered acceleration distributions also retained their shape, indicating preserved local Gaussian equilibria for $t<\tau_{2}$. However, the velocity distributions developed stretched tails, consistent with gradual clustering in the safe cell, and non-gaussianity persisted due to the interplay of $v$-dependent noise and Superstatistics. The only clearly altered observable was $\sigma(v)$, which maintained isotropy but evolved over time. To quantify this evolution, Eq. (4) was fitted independently in successive time windows during the panic regime, yielding one triplet $(\Sigma,\Delta\Sigma,v_{\sigma})$ per window. Figure 4a shows the resulting scalar time series: $\Delta\Sigma(t)$ and $v_{\sigma}(t)$ underwent a sharp, brief increase at minute 9.2, while $\Sigma(t)$ stayed constant within uncertainty.
Figure 4: Temporal Signatures of Panic and Cluster Stabilization in the Panic Regime. (a) Time evolution of the parameters in Eq. (4), obtained from independent fits in successive panic-regime time windows. $\Delta\Sigma(t)$ and $v_{\sigma}(t)$ show a sharp change at minute 9.2, while $\Sigma(t)$ remains stable. (b) Probability of finding an ant from the safe cell at a distance less than or equal to $2\,\mathrm{cm}$ from the barrier. The last 5 minutes of the experiment are highlighted in orange, where the curve remains approximately constant near zero. During this interval, the ant cluster is stable.
Because $\Sigma(t)$ showed no systematic variation, in later simulations it was fixed to the baseline value $\Sigma=0.34\,\mathrm{cm\,s^{-3/2}}$ already reported in Table 1. The remaining parameters were summarized by back-to-back exponential profiles centered at $t_{c}=9.2\,\mathrm{min}$,
$q(t)=q_{0}+A_{q}\,e^{\displaystyle{-\lambda_{q}\lvert t-t_{c}\rvert}},\qquad q\in\{\Delta\Sigma,v_{\sigma}\}$ (9)
where $q_{0}$ is the value away from the peak, $A_{q}$ is the peak excess, and $\lambda_{q}$ is the common growth/relaxation rate. The six scalar coefficients $(q_{0},A_{q},\lambda_{q})$ for $q=\Delta\Sigma$ and $q=v_{\sigma}$ are listed in Table 1 and were used in the later simulations. The peak time $t_{c}$ will subsequently be used as a key measure of one of the most significant timescales in the ants’ mechanism in response to panic.
@@
$\begin{split}\frac{d}{dt}v_{i}(t)&=-\gamma(v_{i}(t)-\tilde{v}_{i}(t))\\ &\quad\,+\sigma(v,t)\sqrt{\hat{D}(t)}\,\xi_{i}(t)+F_{w}(r_{i})\end{split}$ (10a)
@@
V Discussion
The slowly varying diffusivity $D(t)$ does not necessarily correspond to a single ants physiological knob, instead, it is more naturally interpreted as an emergent combination of several factors that jointly regulate ant mobility under confinement. In this context, $D(t)$ can be viewed as the product of several weakly dependent, positive stochastic factors. Taking the logarithm transforms this multiplicative structure into a sum of random contributions, so that $\log D$ becomes the sum of many small terms. Under these conditions, the Central Limit Theorem applies to $\log D$, leading to Gaussian statistics for $\log D$ and, consequently, to Log-normal fluctuations of $D$ itself.
At the biological level, these multiplicative stochastic factors can be naturally associated with the mechanisms by which ants communicate and interact while moving under confinement. Ant-ant communication in Atta insularis likely involves different ways of short-range chemical communication using various parts of the insect’s body [44], and perhaps also substrate vibrations [45]. These would induce a multiplicative contribution to the fluctuating diffusion coefficient: vibrational signals propagate through contact chains, where one information-emitting ant can trigger successive activations in its neighbors, effectively amplifying the signal. Under repeated transmissions, this mechanism is structurally similar to autocatalytic growth, in which successive positive factors are multiplied together [46].
Complementarily, from a coarse-grained kinematics perspective, diffusivity can be understood in terms of effective step lengths and times ($D\sim{\ell_{\text{eff}}^{2}}/{\tau_{\text{eff}}}$), as in generic run-reorient-pause dynamics, where both quantities are themselves composite [47, 48]. Locomotor vigor, intermittency due to pauses, contact rate, or turning dynamics could gate one another multiplicatively, so mobility can be expressed as a product of positive, slowly fluctuating contributions rather than as an additive control parameter [48, 49, 50].
In confinement, effective space available could be a strong multiplicative constraint, simultaneously modulating encounter rates, turning frequency, and the propensity to pause [51, 52, 53, 54]. Small fluctuations in local density can therefore trigger a cascade-like process: changes in density alter collision rates [55], which in turn modify turning and stop-go statistics, feeding back onto local density itself. Such feedback naturally constitutes a multiplicative cascade in space and time, for which Log-normal statistics are a well known outcome [2, 56, 57]. In parallel, ant locomotion can be understood as a stochastic switching among internal movement modes (e.g., resting, slow cruising, fast cruising), with both the residence time in each mode and the transition rates between modes influenced by local encounters [58, 59]. The resulting effective mobility then emerges as a product of locomotor state, time spent in that state, and interaction-mediated gating, again pushing toward Log-normal $D$.
Within this picture, the emergence of Log-normal diffusivity is not driven by transitions between distinct behavioral programs at a colony level (e.g., baseline exploration versus panic response), but by interaction-mediated processes that continuously reshape individual trajectories at the collective level. Panic, therefore, does not create a new dynamical regime; instead, it transiently perturbs a system that is already governed by multiplicative processes. Because the possible multiplicative sources discussed above remain active during panic, the mechanisms responsible for generating Log-normal fluctuations are preserved, endowing the system with a form of statistical resilience. In this sense, confinement could act less as a passive boundary condition and more as a regulator of collective stochasticity.
@@ VI Conclusions
In summary, we have shown that Log-normal Superstatistics emerges as a robust statistical framework that captures the transition from normal to danger behavior in confined ants, while maintaining continuity and providing a foundation upon which panic-induced dynamics can be superimposed. We reported non-Gaussian behavior, a slowly varying diffusion coefficient, and a collective escape response consistent with interaction-mediated information transfer. A data-driven superstatistical model with Log-normal diffusion successfully reproduced these features and captured the danger transition, relying on a single free parameter.
The model uncovers a fast, memory-limited information cascade followed by a slower collective reorganization into a stable cluster, and strongly suggests that the diffusivity itself is not a single physiological control parameter but an emergent quantity arising from the multiplicative combination of interaction-mediated processes under confinement. When expressed as a product of weakly dependent stochastic factors, the logarithm of the diffusivity becomes additive, naturally yielding Log-normal fluctuations via the Central Limit Theorem. The persistence of this statistical structure under panic indicates that the underlying multiplicative mechanisms remain intact, revealing a form of collective statistical resilience and positioning confinement as a regulator of stochastic organization in living active matter.
VII Acknowledgements
@@
A. Reyes gratefully acknowledges the Abdus Salam ICTP STEP program for financial support. We thank A. Celani, D. Gordon, R. Mulet, T. Tohme and D. Machado for valuable discussions, G. Simon for the 3D-printed of the cells, and A. Serrano for assistance with the design of Fig. 1. E. Altshuler acknowledges an Invited Professorship supported by CNRS at the University of Lyon.
@@
Figure 5: Boundary Effects and Effective Wall Thickness. (a) Cartesian acceleration components as functions of Cartesian position, with error bars showing the standard error of the mean. Inset: acceleration absolute value as a function of the distance from the walls. A characteristic length $\ell_{w}=0.6\,\mathrm{cm}$ is defined as the effective wall thickness. (b) The grey sections have been removed from a sample trajectory, as they are within a distance $\ell_{w}$ from the walls: in order to ensure that statistics is not contaminated by boundary effects, we only use for analysis the non-gray sections.
Appendix A: Boundary effects
@@
The baseline regime was used to evaluate boundary effects while measuring the extent of wall “forces”. Experimental Cartesian acceleration components (i.e., $a_{x}$ and $a_{y}$) were arranged based on increasing values of the corresponding Cartesian position components $x$ and $y$, respectively. Both acceleration time series were binned and a running average was taken (main panel of Fig. 5a). The ants slow down as they approach the walls, revealing the influence of the boundaries as if they were confined by a soft, square-box-like potential. Acceleration absolute value against the distance from the wall revealed a characteristic length $\ell_{w}=0.6\,\mathrm{cm}$, obtained by fitting an exponential function (inset of Fig. 5a). Henceforth, analysis was conducted with trajectory segments that were consistently at least $\ell_{w}$ units away from every boundary, to discard boundary effects (Fig. 5b).
@@
Figure 6: Barrier Designs and Contact Pathways for Danger Transmission. (a) Ant used in experiments: leafcutter ant Atta insularis, a species native to Cuba known as “bibijagua”. (b) Sketch of the three types of barriers used for the protocols Glass, Thick and Thin. The latter is also used for the Ctrl protocol. (c) Top view of ants near a thick and thin barriers. Short distance contacts are not allowed by the thick barrier. (d) Lateral view of an ant near a thin barrier. As a reference, the length of the ants bodies is around $1\,\mathrm{cm}$.
Appendix B: Mechanisms for danger transmission
To isolate potential mechanisms of danger transmission, we used three barrier types (Fig. 6b): a transparent Glass wall allowing only visual cues; a Thick barrier restricting the interaction to mutual antennation; and a Thin barrier allowing full contact between ants in adjacent arenas (Fig. 6c).
@@
To characterize collective motion across protocols, we tracked the ants’ CM in both arenas (Fig. 7a-d). In all cases, the CM in the danger cell remained near the geometric center regardless of barrier type. The same held for the safe cell, except under the Thin protocol, where the CM systematically shifted away from the barrier after repellent was introduced (Fig. 7c). Fig. 7e-h shows the time evolution of the horizontal CM position in the safe cell, where mean and standard deviation are shown as a function of time, averaged over 10 experimental repetitions. A clear displacement emerges only in the Thin case after the injection at minute 5. In contrast, the Ctrl case (i.e., where the danger cell is empty) shows no comparable motion, indicating that passive diffusion alone cannot account for the observed shift.
@@
$\Delta\Sigma_{0}$ $A_{\Delta\Sigma}$ $\lambda_{\Delta\Sigma}$ $v_{\sigma_{0}}$ $A_{v_{\sigma}}$ $\lambda_{v_{\sigma}}$
$\mathrm{(cm\,s^{-3/2})}$ $\mathrm{(cm\,s^{-3/2})}$ $\mathrm{(s^{-1})}$ $\mathrm{(cm/s)}$ $\mathrm{(cm/s)}$ $\mathrm{(s^{-1})}$
1.53 3.07 0.09 1.25 5.48 0.06
Figure 7: Barrier-Dependent Panic Transmission and Center of Mass Analysis. Snapshots of experiments (a) Glass, (b) Thick, (c) Thin and (d) Ctrl. Green, blue, and red traces show the center of mass (CM) trajectory for consecutive 5-minute windows. (e-h) Horizontal CM evolution in the safe cell, relative to its initial position. Shading shows standard deviation. Repellent is added to the danger cell at $t=5$ min. (i) Net horizontal displacement of the ants CM based on 10 repetitions of each type of experiment. Mann-Whitney tests assert to the Thin protocol to statistically differs from the other three.
Figure 8: Persistence of Superstatistical Signatures Under Panic. Panels (a-f) were computed over consecutive time segments. The first segment corresponds to the baseline interval $[0,5)$ min, whereas the panic regime was analyzed using shorter windows; the legend reports the final time of each segment, so $[14,15)$ denotes the last 11-min interval. (a) Speed dependence of parallel and perpendicular acceleration components across time intervals. (b) Local flatness of Cartesian velocity components vs lag time, computed over those trajectory segments. (c) Velocity autocorrelation functions (VACF) vs lag time. The solid line has slope $\gamma=(0.53\,\mathrm{s})^{-1}$. Notice that the slow time scale $\tau_{2}$ is not captured by the VACF. (d) Probability density function (PDF) of centered parallel and perpendicular components of acceleration fluctuation. (e) PDF of Cartesian velocity components. (f) Temporal and speed dependence of the intensity of acceleration fluctuations for the parallel (blue) and perpendicular (orange) components after panic was induced. Continuous lines show fits to Eq. (4).